OpenStudy (anonymous):

(-1/64)^(-1/3) is equals to -4 right?

hartnn (hartnn):

(-1/64)^(-1/3) is a complex number, are you studying complex numbers topic?

OpenStudy (anonymous):

we consider \[i=\sqrt{-1}\] in case of sq. root but what about\[\sqrt[3]{-1}\] in case of cube root? what we call it as? @hartnn

hartnn (hartnn):

there is no specific name for that as such

OpenStudy (anonymous):

it says to evaluate the expression. Is it -4 the ans?

hartnn (hartnn):

do you have the options/choices?

OpenStudy (anonymous):

no the ans. should contain \[\sqrt[3]{-1}\]

OpenStudy (anonymous):

a. -1/4 b. 4 c.262,144 d. -4

hartnn (hartnn):

*sigh* then they are considering \((-1)^{2/3} = 1\) which is not true ... from the options, i would go with -4 only...

OpenStudy (anonymous):

ans. may be \[4*\sqrt[3]{-1}\]

hartnn (hartnn):

\(-4 (-1)^{2/3}\) is the actual answer :) or \(4\angle 60\)

hartnn (hartnn):

oops \(4 \angle -60 \)

OpenStudy (anonymous):

i think\[\sqrt[3]{-1}=\sqrt[3]{(-1)(-1)(-1)}=-1\] if this is correct, the answer is -4. @hartnn @bless

OpenStudy (anonymous):

please see to it

hartnn (hartnn):

cube root of -1 is again complex :) \(1\angle 60\)